6.3 (LO 1),AP Barnes Company reports the following operating results for the month of August: sales 000 (units 5,000); variable costs $210,000; and fixed costs $75,000. Management is considering the following independent courses of action to increase net income. I. Increase selling price by 10% with no change in total variable costs or sales volume, 2. Reduce variable costs to 58% of sales. 3. Reduce fixed costs by $15,000 Instructions Compute the net income to be earned under each alternative. Which course of action will produce the highest net income?
Answer #1

Answer

  • All working forms part of the answer
  • Net Income under each alternative

Current

Alternative 1

Alternative 2

Alternative 3

Sales

$325,000

$357,500

$325,000

$325,000

Variable Cost

$210,000

$231,000

$188,500

$210,000

Contribution margin

$115,000

$126,500

$136,500

$115,000

Fixed Cost

$75,000

$75,000

$75,000

$60,000

Net Income

$40,000

$51,500

$61,500

$55,000

--Working for above

Current

Alternative 1

Alternative 2

Alternative 3

Sales

325000

=325000 + (325000 x 10%)

325000

325000

Variable Cost

210000

=210000*357500/325000

=325000*58%

210000

Contribution margin

=325000-210000

=357500-231000

=325000-188500

=325000-210000

Fixed Cost

75000

75000

75000

=75000-15000

Net Income

=115000-75000

=126500-75000

=136500-75000

=115000-60000

  • Highest Net Income would be in Alternative #2: reducing Variable cost to 58% of sales.